Suppose a new car is purchased for $46,357 and depreciates by 21% over the first year of ownership. If the car is driven 12,570 miles in that year, what is the cost, in dollars, per 100 miles driven, for depreciation?
Answer #1
Depreciation for first year = Cost * 21% = 46357 * 21% 9734.97
Depreciation per 100 miles = ( Depreciation for first year / Miles driven in first year ) * 100 = ( 9734.97 / 12570 ) * 100 77.45